OpenStudy (anonymous):

solve for x. x - 8x + 7 = 2x - 11

OpenStudy (grant330sims):

first, try to get all the variables on the same side.

Parth (parthkohli):

-7x + 7 = 2x - 11 => -7x + 18 = 2x => 7x + 2x = 18 => 9x = 18 => x = 2

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

x - 8x + 7 = 2x - 11 steps: -7x+7=2x-11 -7x-2x=-11-7 -9x=-18 x=-18/-9 x=2